# SMILES has to go : Representation of Molecules via Algebraic Data Types

### v1

@@ Abstract
This paper proposes a novel representation of molecules through Algebraic Data Types (ADTs). The representation has useful properties primarily by including type information. The representation uses the Dietz representation enabling representation of organometallics with multi-centre, multi-atom bonding and delocalised electrons, resonant structures and co-ordinate data of atoms. Furthermore, this representation goes further than any other in the literature, providing a natural data structure to represent shells, subshells and orbitals. Perks of the representation include it’s natural inclusion in reaction descriptions and the ability to make molecules instances of algebraic groups. The representation is further motivated as providing guarantees for those wishing to do Bayesian machine learning (probabilistic programming) over molecular structures. A criticism of competing and commonly used representations such as SMILES and SELFIES is provided and solutions are proposed to the weaknesses of these along with an open source library, written in Haskell. An example of integrating the library with LazyPPL – a lazy probabilistic programming library written in Haskell – is provided, conceptually justifying the efficiency of the representation over string based representations and recent work such as SELFIES [Krenn et al., 2020]. This library distinguishes between the data and the type of data – enabling a separation of concerns between interface and object. I solve three problems associated with the future of SELFIES, molecular programming language, 3D information, syntactic invalidity and Dietz representation.
Keywords: Molecular Representation, Cheminformatics, Algebraic Data Types
1Oxford University, Department of Computer Science, oliver.goldstein@cs.ox.ac.uk
1 Scientific Contribution
@@
    Introduction of a minimal grammatical representation of molecules. This is combined with detailed criticism of previous molecular representations, along with proposed solutions for each issue identified. Perks of the representation are explored including naturally modelling reactions and group theoretic properties of molecules.
@@
    Conceptual justification of efficient Bayesian Inference over this representation demonstrated by an example with a Haskell implementation using lazy trace Metropolis Hastings and 3D molecular generation [Paquet and Staton, 2021].
2 Introduction
The choice of molecular representation is consequential for the cheminformatician, and there is an ongoing search for computational representations which produce superior results in machine learning (ML) tasks such as novel molecule discovery, molecular property prediction, and reaction predictions. Usable computational representations which are capable of representing certain types of molecules, or capable of adequately representing reactions in a computational format, are still being searched for. Currently, string based representations are the predominant type of representation used and studied by cheminformatics for such tasks, but they are not without problems.
SMILES [Weininger, 1988] is the most commonly used string based representation, particularly in the case of machine learning [Gómez-Bombarelli et al., 2018], [Zheng et al., 2019] & [Deng et al., 2022]. Its well-known problems include the fact that not all valid SMILES strings are valid molecules. SMILES strings also cannot represent complicated bonding systems (such as in organometallics), resonance structures, any kind of delocalised bonding, and certain kinds of complex stereoisomerism, problems shared with other representations which use molecular graphs. In addition, small changes in SMILES strings can lead to large changes in a molecule’s structure, and adjacent atoms in the molecule are often far apart in the string making mathematical measures of similarity difficult to calculate [Fang et al., 2022],
Different fixes for the problems of string based representations have been attempted or proposed, but these usually require the creation of brand new grammars, of which there is no shortage.
In this paper, we summarise and critique modern computational representations, and proposed representations, specifically with a focus for ML applications. In particular, we critique string based representations and their numerous extensions, including some limitations which are under-discussed by the cheminformatics community, such as the limitations of string based representations in the application of probabilistic programming.
Then, as a solution to these identified problems, we introduce a novel representation of molecules: the Algebraic Data Type (ADT).
Our ADT is written in the pure, type-safe, functional programming language Haskell, and implements a modified Dietz representation, with additional (and optional) molecule properties such as 3D coordinate information and quantum information. It easily represents resonance structures, delocalised bonds and organometallics, and due to the inclusion of quantum information, can be used for quantum chemistry applications.
We also demonstrate how the ADT can be easily extended to represent reaction networks, and we provide a simple example of a probabilistic model which generates molecules and scores them with an arbitrary distance metric for a chosen molecule.
Despite the immediate benefits presented by our ADT, this paper does not does not intend to provide an actual implementation of a sound and complete molecular representation. It merely suggests ADTs as a form of representation, and instead, emphasis is placed on the advantages of ADTs and its potential extensibility and use with probabilistic programming. We leave it to the cheminformatics community to discover the ways in which ADTs would be most useful.
2.1 Types of representation
This paper distinguishes between different usages of ‘molecular representation’.

### v2

# SMILES & SELFIES have to go: Representation of Molecules via Algebraic Data Types

@@ Abstract
The Algebraic Data Type (ADT) can be used as a computational framework for molecular representation for the purpose of advancing tasks in cheminformatics. This can include generative modles in the context of Bayesian machine learning via probabilistic programming. The ADT that we put forward, implements the ‘Dietz’ representation for molecular constitution via multigraphs of electron valence information, and uses 3D coordinate data to provide stereochemical information, easily enabling the representation of complex molecular phenomena such as organometallics, multi-center bonds, delocalized electrons, and resonant structures. Unlike traditional string-based representations such as SMILES and SELFIES, the ADT is much more flexible, yet retains desirable qualities from type-safety to seamless integration with Bayesian Probabilistic Programming. An extensive criticism of both SMILES and SELFIES, in this article, is given, along with criticisms of the so-called Future of SELFIES. An open-source library implemented in Haskell demonstrates the ADT along with experimental extensions demonstrating its use in reaction modelling, group theoretic applications, and integration with LazyPPL, a lazy probabilistic programming library. Also provided as an extension is the ability to represent electronic structures, including shells, subshells, and orbitals, significantly expanding its representational scope compared to other representations in the literature. These features position the proposed ADT as a robust alternative to existing molecular representations, addressing limitations such as inadequate support for 3D information and syntactic invalidity while offering a platform for innovative cheminformatics research. Accompanying discussions about the meaning of a ‘representation’ are included. The fully functioning GitHub library can be found at https://github.com/oliverjgoldstein/chemalgprog.
keywords:
Molecular Representation, Cheminformatics, Algebraic Data Types
1 Introduction
The choice of molecular representation directly affects cheminformatics tasks such as novel molecule discovery and molecular property prediction [1]. More expressive representation enables the exploration of larger regions of chemical space, both as input to and output from machine learning models. However, representational ability alone does not determine the effectiveness of a molecular representation in ML applications. Computational efficiency of optimisation algorithms, model choice and the amount of data provided are also important factors, along with the ease—or even possibility of integrating representations with techniques such as geometric deep learning [2] or Bayesian machine learning [3]. In cheminformatics, we believe that when it comes to molecular representation, “the medium is the message.”.  [^1] From Marshall McLuhan. More specifically, the context of communication decides the operations which the transported data supports.
Despite decades of research, string-based representations such as SMILES [4] and SELFIES [5] remain the most common representations for ML tasks [6, 7, 8], despite well documented short-comings in representational ability and with syntaxes themselves. For instance, standard SMILES strings cannot represent complex bonding systems, resonance structures, or delocalized bonding, and do not include any 3D information, with only a limited ability to describe stereochemical features. In addition, not all SMILES strings correspond to valid molecules, meaning that the output of ML models is, in many cases, gibberish [5], and small changes in SMILES strings can result in large differences in molecular structure, making property prediction difficult. Attempts to build upon strings have led to a profusion of new grammars and syntaxes to represent molecular phenomena, a problem in itself, such as BIGSMILES or CURLYSMILES etc.
While proposing future work to address the limitations of string-based representations such as SELFIES, [5] suggest developing a “molecular programming language”. Reasoning that strings can encode Turing-complete programming languages (e.g., Python source code), they suggest SELFIES could itself become a programming language, extended with a syntax that allows for recursion (for/while loops) and logical statements. Although this reasoning appears to conflate strings as a data structure with strings as a sequence of ASCII characters, we ask the question: if molecules can be represented as data within well established programming languages, we ask, why not simply use one?
In this paper, we critique contemporary computational representations, their application and challenges with ML, and some of the proposed solutions relating to string-based representations — we then present a solution with a novel framework: the Algebraic Data Type (ADT), written in the functional and typed language, Haskell.
The ADT has features of the Dietz representation for constitution [9], using multigraphs of atomic valence information. It contains optional features such as 3D atomic coordinate data to provide functions such as conformation energy prediction and configuration of molecules, as well as quantum information in the form of orbitals, which is a first in the literature.
The ADT is able to represent complex bonding systems (e.g. organometallics), resonance structures, and delocalised bonding. We further demonstrate extending the ADT in modelling reaction networks, and by using the ADT as an instantiation of a mathematical group. We also construct a probabilistic model in LazyPPL that generates and scores molecules based on arbitrary metrics to show the ADT’s ability to be used with probabilistic programming.
We also consider the impact the ADT has beyond computation, and consider the impact of the ADT as a representation for storage (and retrieval) of molecular information, and for transmission of molecular information between cheminformaticians.
While not intended as a definitive implementation of molecular representation, the ADT demonstrates the potential of algebraic approaches to reshape computational chemistry and invites the cheminformatics community to explore its capabilities and extend its utility.
2 Forms of Molecular Representation
In this paper, we distinguish three different usages of the term “molecular representation”:
@@ 2 Forms of Molecular Representation
    Storage representations refer to the format for storage of molecular data on-disk, that is, without interpretation by a computer program.
@@
    Transmission representations relate to the way molecules are sent between cheminformaticians, and may depend on the medium of communication used.
  3. 3.
    Computational representations refer to how molecules are represented inside programming language or program, specifically, the data type or types which are used to represent the program.
For instance, a molecule may be stored in a database as a molfile, be processed with cheminformatics library such as RDKit[10] into another internal representation, and be then transformed into a molecular graph, or a fingerprint, for processing in a machine learning model[11, 12].
String based representations, such as the ubiquitous SMILES[4], may be used for all three. They are available in databases, may be downloaded and transmitted as strings, and also used as direct input for machine learning models[11, 12, 1], which as we later discuss, we regard as a mistake.
We regard the distinction between underlying data, and the data type to be crucial in cheminformatics, as it is the data type which constrains the operations which may be performed on data, reflects the structure and semantics of the data, and affects the efficiency and utility of computational tasks. Failure to make this distinction can lead ambiguity, inconsistency, and potential errors in interpretation and manipulation for meaning, as we later discuss in detail.
3 Contemporary Molecular Representations
In this paper we consider Molecular Graphs, Molecular Hypergraphs, the Dietz representaion (via multigraphs), string based representations and their numerous extensions, fingerprints, Variational Auto Encoders (VAEs), and computation tables.
3.1 Molecular Graphs
Many common molecular representations are built around molecular graphs and, as such, inherit strengths and limitations of molecular graphs for molecular representation [11]. String based representations such as SMILES and SELFIES use a number of rules to decode ASCII strings which correspond to molecular graphs, and MDF molfiles have “blocks” which correspond to sets of atoms and sets of bonds.
Definition 1 (Graph).
Formally, in set-theoretic notation, a graph, $G=(V,E)$ is a pair of sets $V$ and $E$, where $V$ is a set of vertices and $E$ is a set of pairs of vertices {u,v}, called edges.
A molecular graph is an instance of a graph in which the vertices are atoms, and the edges are bonds between atoms. Molecular graphs are connected, undirected, and labeled, where vertices may be labeled with information such as chemical element, and charge; and edges may be labeled with bond type (e.g. single, double or aromatic.), bond order, or 3D information (e.g. R/S isometry) [9, 11].
One benefit of representing molecules computationally as graphs is that insights from chemical graph theory can be directly leveraged by machine learning models. However, not all molecules can be represented by graph models, including molecules with complex bonding, organometallics, resonant structures and tautomerism.
Edges in graphs are between two vertices only, so molecular graphs have a limited ability to represent delocalised and multi-centre bonding, including ionic bonds and organometalics [11, 12]. Although it is possible to label edges with, e.g. rational bond orders, the contribution of individual electrons to multiple covalent bonds would not be described by such a graph. Attempting to express delocalised bonding by labeling edges of a molecular graph with bond types would require a sufficient number of predefined bond types, which would need to be extended for each new bonding relation not described.
Evidence of this approach extends back decades, as Dietz said ([9]):
“Enhancing the expressiveness by including a new bond type for every exceptional case is certainly not a very elegant solution.”
Using a single structure, graph models cannot represent tautomers, so instead require a work-around such as using multiple structures to represent the different forms of molecules.
Hypergraph and multi-graph models have been proposed to solve some issues with graph models [13], but they have seen limited use [11].
3.2 Molecular Hypergraphs
A hypergraph, is a graph where edges may connect any number of vertices. In a molecular hypergraph, edges with two vertices represent localised bonding, and edges with more than two verticies represent delocalised bonding, and are labelled with the number of electrons shared between the corresponding atoms.
From Dietz:
“A hyperedge containing more than two atoms gives us no information about the binary neighborhood relationships between them. That means we have no information at our disposal conceming the way in which the electrons are delocalized over these atoms” [9].
Hypergraphs would indicate an electronic “soup” with out knowledge of the manner electrons are delocalised. Hypergraphs remain uncommon.
3.3 Multigraphs
Multigraphs are not widely used as a standard computational representation in cheminformatics or chemical graph theory, despite the protestations of Dietz [9]. Existing tools and frameworks generally rely on labeled graphs or weighted graphs to encode bond multiplicities and types.
Theoretical discussions of multigraphs in chemical graph theory arise in the context of molecules with multiple bond types or delocalized bonding, but computational implementations of multigraph representations are not widespread.
Definition 2 (Multi-Set).
A multiset (or bag) is a generalization of a set where elements are allowed to appear more than once. Formally, a multiset $M$ over a universe $U$ is defined as a function:
$M:U\to\mathbb{N}$
where $M(x)$ denotes the multiplicity of element $x$, i.e., the number of times $x$ appears in the multiset. A finite multiset can be represented as:
$M=\{a_{1}^{m_{1}},a_{2}^{m_{2}},\dots,a_{n}^{m_{n}}\}$
where each element $a_{i}$ appears $m_{i}$ times.
The cardinality of a multiset $M$ is given by:
$|M|=\sum_{x\in U}M(x)$
which counts the total number of elements, including repetitions.
Definition 3 (Multi-Graph).
Formally, in set-theoretic notation, a multi-graph $G=(V,E)$ is an ordered pair of a set of vertices, $V$, and a multi-set, $E$, of edges.
By using a multigraph approach, it is possible to label bonds with information about the binary bonding relationships between atoms, as well as describing the delocalised electrons that persist over a subset of atoms, which is the approach taken by Dietz [9].
3.4 Dietz representation
The Dietz representation was designed to provide a computational framework capable of representing any chemical structure that can be described by a structural formula. It was developed to overcome the limitations of graph-based and hypergraph-based methods, including string representations that correspond to molecular graphs [9].
The paper describes the “constitution” of a structure in terms of the unshared valence electrons of atoms and the bonding relationships between them. Additionally, the terms “conformation” and “configuration” are used to describe different aspects of the relative spatial relationships between atoms. Conformation refers to the spatial arrangement of atoms that can change without breaking chemical bonds, such as rotations around single bonds, while configuration refers to the fixed spatial arrangement of atoms, including stereochemistry, which requires breaking and reforming bonds to alter. These spatial relationships are illustrated using a unique “paddle wheel” representation.
As the proposed ADT uses atomic coordinate data to describe 3D information of molecules, only the “constitution” of the Dietz representation is discussed here.
The constitution of a structure is represented in set theoretic notation, with two sets. Our ADT includes type information, which prevents invalid molecules as will be discussed. The first set contains all atoms in the structure as triples, containing the number of unshared valence electrons of the atom, a numerical index, and the atomic symbol. The second set describes “bonding systems”, or bonding relationships between atoms, and contains pairs consisting of the number of shared valance electrons, and a set of edges representing bonds over which the electrons are shared.
Definition 4 (Constitution).
The constitution of a molecule is an ordered pair:
$C=(V,B),$
where $V$ is the set of vertices, representing the atoms in the molecule, and $B$ is the set of bonding systems, representing the bonding relationships between atoms.
Each vertex in V is a triple:
$V=\{(u,j,A)\mid u\in\mathbb{N}_{0},\,j\in\mathbb{N},\,A\in\text{Symbols}\},$
where:
  • u is the number of unshared valence electrons for the atom,
  • j is a unique index, and
  • A is the atomic symbol (e.g., $H,C,O$).
Each bonding system in B is a pair:
$B=\{(s,\{\{x,y\}\mid x,y\in V\})\mid s\in\mathbb{N}\},$
where:
  • s is the number of shared electrons in the bonding system, and
  • {{x, y}} is a set of unordered pairs of vertices representing bonds over which the electrons are shared.
Localised bonding is represented by a bonding relationship containing a single set of edges. For instance, a single covalent bond between two atoms would be represented as follows: $(2,\{\{a_{1},a_{2}\}\})$, where $a_{i}$ is the index of each atom.
Delocalised bonding is represented by a bonding relationship containing multiple set of edges. For instance, the delocalised pi electrons of a benzene ring would be described by a bonding relationship of the form $(6,\{\{c_{1},c_{2}\},\{c_{2},c_{3}\},\{c_{3},c_{4}\},\{c_{4},c_{5}\},\{c_{5},c_{6}\},\{c_{6},c_{1}\}\})$, where $c_{i}$ is the numerical index of each carbon atom in the ring.
This representation is equivalent to a multigraph with additional labelling or metadata to describe the groups of edges with shared electrons.
Example 1 (Constitution of Hydrogen).
The constitution of hydrogen
$V(H_{2})=\{(0,1,H),(0,2,H)\},$
$B(H_{2})=\{(2,\{\{1,2\}\})\}\vskip 6.0pt plus 2.0pt minus 2.0pt$
There are two hydrogen atoms, labelled 1 and 2, each with zero unshared valance electrons in the atomic core, and a single covalent bond with two shared valance electrons.
Example 2 (Constitution of Benzene).
The constitution of Benzene:
$V(Benzene)$ =$=\{(0,1,H),(0,2,H),(0,3,H),(0,4,H),(0,5,H),(0,6,H),$
[…]
The delocalised pi electrons of the benzene ring are described in a single bonding system, as the final element of the set, and separately from all other covalent bonds in the structure.
A major advantage of the Dietz representation is its general model for the representation of bonding. As a result, it is straightforward to describe complex bonding systems such as organometallics, with ferrocene and diborane being given as examples in the original paper. The lack of labeling requirements for delocalised bonding stands in contrast to string based representations such as SMILES, which has explicit rules in the grammar (a form of labeling) for aromatic carbon, and two rules for different cases of aromatic nitrogen.
Despite satisfying the three requirements identified in [1] for representation beyond organic chemistry—namely, the ability to represent delocalized bonding, explicit labeling of terminal hydrogen atoms, and advanced stereochemistry such as in cumulenes—Krenn et al. describe the Dietz representation (apparently mischaracterized as using hypergraphs) as one that “leads to complicated nested sets of brackets that may be hard to comprehend” [1].
Although set-theoretic notation may be unfamiliar, the definition or “grammar” of the constitution of a molecule can be efficiently described. This stands in contrast to SMILES strings or SELFIES strings (which as an additional challenge contains overloading of symbols in the grammar [1]).
As Krenn [1] puts it:
The same applies to other weak interactions such as hydrogen bonds, raising important questions as to which interactions should actually be captured in a digital molecular representation as a “bond” (and which should not) and how to automatically detect them from a set of atomic coordinates, ultimately leading to a rather arbitrary distinction between bonded and non-bonded. To quote Democritus: “Nothing exists except atoms and empty space; everything else is only opinion”.
3.5 SMILES
SMILES (Simplified Molecular Input Line Entry System, [4]) encodes molecular graphs as ASCII strings through a series of rules, representing atoms and bonds as characters and symbols. Introduced in the 1980s, SMILES was designed to facilitate chemical database searching and to enhance portability, and has achieved significant popularity in cheminformatics [11, 12, 1, 14, 15, 16, 17]. However, increasing demands of modern machine learning and cheminformatics reveal challenges that extend beyond the scope of SMILES’ original design.
As molecular graphs, SMILES inherit the limitations of this representational framework. While SMILES can represent aromaticity through specific syntax[4], it cannot encode delocalized bonding beyond such cases. Resonance and tautomerism must either be simplified into a single structure or require multiple distinct structures to represent each form.
The encoding of stereochemistry in SMILES has notable constraints. While tetrahedral centers are well-supported through isomeric SMILES, a common extension, the encoding of more complex stereochemical features, such as dynamic or rotational stereoisomerism seen in atropisomers, remains unsupported [18]. Isomeric SMILES can represent certain forms of axial chirality in cumulenes and handle non-tetrahedral stereocenters, including allene-like, trigonal-bipyramidal, and octahedral configurations [18]. However, these representations rely on 2D graph-based encoding and lack support for nuanced 3D spatial arrangements required for complex organometallic systems or multi-center bonding scenarios. The absence of three-dimensional (3D) spatial data further limits SMILES’ utility for applications requiring dihedral angles or precise molecular conformations, a critical aspect of protein-ligand modeling and other machine learning tasks such as property prediction [19, 20, 21].
Beyond these limitations, SMILES suffers from unique structural and functional shortcomings such as syntactic invalidity [5, 11, 12], non-local encoding of features, and non-unique representations of molecules, all of which can impair machine learning models as the models must implicitly learn the SMILES syntax, which can be difficult, and require larger training sets and wasted computation [11, 22].
“Syntactic invalidity” means that not all valid SMILES strings correspond to valid molecules. In Machine learning contexts, algorithms may output syntactic nonsense, and must re-sample or require larger training sets to ensure robustness [11, 6, 5]. Even though one paper claims that invalid SMILES should be seen as beneficial to chemical language models [23], it should be noted that the perceived improvement is only relative to comparisons between SMILES and SELFIES models, and likely reflects the models’ enhanced ability to infer SMILES syntax when generating valid molecules.
In SMILES, small changes in the string can lead to drastic changes in molecular structure, and vice versa – adjacent atoms in the molecule can be far apart in the string [22]. As well as impacting readability, the non-local encoding of structural information can make it difficult to write or even find a similarity metric of SMILES strings whose magnitude corresponds to a relevant similarity of molecules [22].
SMILES representations are non-unique: a single molecule can be expressed by multiple valid strings [4]. As well as complicating database searches [12], machine learning algorithms must discern the equality of different representations. Although some machine leaning approaches have attempted to utilize multiple SMILES for the same molecule in their training sets, and have claimed improvements to their models as a result [24, 25], this is again likely explained by the ability of the model to infer the SMILES syntax more easily. In other words, it exposes a weakness of SMILES for machine learning, not an advantage.
Representations in SMILES also do not give indications as to molecular symmetry. By example, the Cyclopentadienyl anion in SMILES is [cH-]1cccc1, but this does not indicate the rotational symmetry of the molecule, as is clearly shown in Fig. 1. The non-uniqueness of SMILES strings can be seen as 1cccc1[cH-] is also a valid representation, but in SMILES no explicit indication is given as to whether [cH-]1cccc1 = 1cccc1[cH-].
Figure 1: The Cyclopentadienyl anion. Although the structure displays clear rotaional symmetry, this is not indicated by the SMILES representaion: [cH-]1cccc1. Image Credit: Oliver Goldstein (Author).
Attempts to canonicalize SMILES, and the proprietary nature of the Daylight canonicalisation algorithm, has led to inconsistencies between implementations and research teams [11], though the OpenSMILES initiative has mitigated this issue by providing a standardized grammar.
Importantly, as a data type, strings by themselves provide no guarantees about the structure of the validity of the data.
While SMILES and its extensions remains useful tools for encoding and sharing molecular structures, its limitations constrain its applicability in modern cheminformatics and machine learning workflows. Extensions to SMILES to fix problems, and the problems they themselves introduce, are discussed in a later section.
3.6 SELFIES
SELFIES (Self-Referencing Embedded Strings) [5] solves arguably the most pressing issue with SMILES: syntactic invalidity. They create a “100% robust” representation, which would translate to soundness in the terminology of logic or computer science. In particular, they present a Context Free Grammar (CFG) where all sequences of terminal symbols represent molecular graphs (soundness) and all molecular graphs can be represented (completeness – but only up to the limited CFG they have created, see Fig. 1). The CFG of SELFIES can be translated into sum and product notation as every CFG can be represented by an algebraic data type [26].
In attempting to retain semantic as well as syntactic validity after mutation, SELFIES uses a table of derivation rules, with overloading of terms to encode valence information [1]. Although successful in maintaining semantic validity after mutation (with the given set of rules, for small biomolecules), and as sucessfully demonstrated with GANs and VAEs[5], the overloading of terms makes the grammar more complicated. The authors claim that SELFIES may be more readable then SMILES for large molecules, although they acknowlede that “Read-ability is in the eye of the beholder”[5]. We personally behold SELFIES as not being human readable, but invite readers to draw their own conclusions.
Machine learning algorithms would also have to implicitly learn the rules of the grammar [11]. Despite the authors claim that SELFIES may be more “machine readable” this remains a conjecture.
The overloading of symbols also represents a challenge for readability. When discussing previous line notations in [4], Weininger criticized attempts to save space with overloaded terms: “Many of the pitfalls in other line notations can be attributed to overuse of symbols and hierarchical rules based on the length of the final notation. Advances in computer hardware have made these restrictions obsolete.” In this sense, SELFIES represents a retrograde step.
State $\epsilon$ [F] [=O] [#N] [O] [N] [=N] [C] [=C] [#C] [Branch1] [Branch2] [Branch3] [Ring]
$X_{0}$ $X_{0}$ F 0 $X_{2}$ N $X_{3}$ 0 $X_{2}$ $NX_{3}$ $CX_{4}$ $CX_{4}$ $CX_{4}$ ign $X_{0}$ ign $X_{0}$
$X_{1}$ $\epsilon$ F 0 N 0 $X_{1}$ $NX_{2}$ $NX_{2}$ $CX_{3}$ $CX_{3}$ ign $X_{1}$ ign $X_{1}$ R(N) R(N)
$X_{2}$ $\epsilon$ F =O $NX_{2}$ 0 $X_{1}$ $NX_{2}$ =N $X_{1}$ =C $X_{2}$ =C $X_{2}$ B(N, $X_{5}$) $X_{1}$ B(N, $X_{5}$) $X_{1}$ R(N) $X_{1}$ R(N) $X_{1}$
$X_{3}$ $\epsilon$ F =O #N 0 $X_{1}$ $NX_{2}$ =N $X_{1}$ =C $X_{2}$C $X_{1}$ B(N, $X_{5}$) $X_{2}$ B(N, $X_{6}$) $X_{1}$ R(N) $X_{2}$ R(N)$X_{2}$
$X_{4}$ $\epsilon$ F =O #N 0 $X_{1}$ $NX_{2}$ =N $X_{1}$ =C $X_{2}$C $X_{1}$ B(N, $X_{5}$) $X_{3}$ B(N, $X_{7}$) $X_{1}$ R(N) $X_{3}$ R(N)$X_{3}$
$X_{5}$ C F 0 N 0 $X_{1}$ $NX_{2}$ $NX_{2}$ $CX_{3}$ $CX_{3}$ $X_{5}$ $X_{5}$ $X_{5}$ $X_{5}$
$X_{6}$ C F =O $NX_{2}$ 0 $X_{1}$ =N $X_{1}$ =C $X_{2}$ =C $X_{2}$ $X_{6}$ $X_{6}$ $X_{6}$ $X_{6}$ $X_{6}$
$X_{7}$ C F =O #N 0 $X_{1}$ =N $X_{1}$ =C $X_{2}$C $X_{1}$ $X_{7}$ $X_{7}$ $X_{7}$ $X_{7}$ $X_{7}$
$N$ 1 2 3 4 5 6 7 8 9 10 11 12 13 14
Table 1: Derivation rules of SELFIES for molecules in the QM9 dataset. Table formatted to fit within the page.
Again, like SMILES, resonant structures are not directly expressible in SELFIES, and delocalized bonding (other than aromaticity) cannot be expressed, along with tautomers.
Since publication, SELFIES has added the ability to represent chirality, through importation of symbols from SMILES, although some forms produce errors in decoding. However, three dimensional information is not included, and it is unclear how to extend the grammar whilst retaining it’s simplicity and soundness. The spatial arrangement of ligands in square-planar stereochemistry, like that of cisplatin and transplatin, cannot be fully captured without 3D information, and so in exploring the chemical space, SELFIES remains unable to represent potentially crucial information for novel drug discover and property prediction, despite being sound.
Furthermore, E-Z stereochemistry cannot provide accurate dihedral or torsional angles [27]. Proteins rely on these angles for many tasks, which may be critical for machine learning tasks [28]. As Dayalan explains ([21]) “Dihedral angles are of considerable importance in protein structure prediction as they define the backbone of a protein, which together with side chains define the entire protein conformation.”. Angles influence the energy associated with a conformation of a drug which in turn relates to the utility and stability of the drug [29]. Although SELFIES solves the syntactic invalidity problem of SMILES, many challenges remain.
3.6.1 The future of SELFIES and of molecular representation - Extensions of SELFIES and future projects
In “SELFIES and the Future of Molecular String Representations,” Krenn et al. (2022) [1], review the limitations of string-based molecular representations in cheminformatics and propose future directions, primarily centered around extending string-based methods.
They highlight several representational difficulties with current molecular string formats, including their inability to effectively capture macromolecular structures, crystal lattices, complex bonding (e.g., organometallics), advanced stereochemistry, precise conformations, and non-covalent interactions such as hydrogen and ionic bonding. These limitations, they argue, are not only present in SMILES and SELFIES but are features of most modern, digital molecular representations in use.
The authors also identify challenges specific to machine learning applications, such as issues arising from the non-local encoding of molecular features within strings, and the impact this has on the latent space of VAEs, as well as the impact overloading of symbols has the efficiency of generative models, as well as human readability.
3.7 Macromolecules and crystals
For macromolecules, Krenn et al. discuss existing approaches such as CurlySMILES, BigSMILES, and HELM, each of which introduces distinct syntactic modifications to account (or not, as the case may be) for repeating subunits, complex connectivity, and stochastic relationships between subunits. They propose BigSELFIES and HELM-SELFIES as future extensions to enable the use of generative models which take strings as input for macromolecular synthesis and design.
In the case of crystals, the authors propose yet another syntax, Crystal-SELFIES, based on the labeled quotient graph of the crystal structure. The formalism of the approach is largely unexplored in the paper, however relies on representation of subunits, in a related manner as with macromolecules.
3.8 Barriers to a general, digital, molecular representation
The representational limitations of current string-based methods are said to rely on three core assumptions: